Solve the system by addition.
x + y = 8
x – y = 8

Answer by ramkikk66(644)

Solve the system by addition.
x + y = 8
x – y = 8
Add the 2 equations to get:
x + x + y - y = 16
i.e. 2x = 16
x = 16/2 = 8
Substituting for x in one of the equations, we get
8 + y = 8, or y = 0.
Solution: x = 8, y = 0
:)